Consider the following static (closed-economy) version of the Classical model:

Y = F (K, L)

C = A + a(Y − T ), with A > 0 and 0 < a < 1,

I = B − br, with B, b > 0,

where A and B represent respectively the autonomous components of consumption (C) and investment (I). Assume the factor inputs, K (capital) and L (labor), are fixed in supply. Finally, assume that government expenditures (G) and taxes (T) are exogenously given. [50 points; 25 each]

(a) Predict and explain the effects of an exogenous increase in government spending (dG > 0) on the interest rate (r), equilibrium output (Y ) and the composition of output.

(b) Now suppose there exists a balanced-budget rule that requires T= G. This requirement, in turn, implies that dT = dG. Predict and explain how does this balanced budget requirement affects the predictions you made above in part (a) regarding the effects of an exogenous increase in government expenditures (dG > 0) on the interest rate (r), equilibrium output (Y ) and the composition of output.
Answer #1

. Ina closed economy, agg demand is equal to the agg supply they ys - yd = c+I+G G = Exogenous value. C = At aly-T) AYO ; oca

let the govt increases tax by the same amt as increase in the goxt spending this dt = dG Changein LT XO! 12 by AY = - MPC - M